You are dragging a heavy box with a mass of 90.0 kg across the floor at a constant velocity with a horizontal force of 300.0 N. a) Calculate uk. b) A friend tries to assist you by pushing up on the box with a force of 100.0 N . How much less force will you need to exert to keep the box moving at a constant velocity? 


1
Expert's answer
2021-01-10T18:30:00-0500

(a)


"\mu_kN=300\to \mu_kmg=300\to \mu_k=300/(90\cdot9.8)=0.34" . Answer


(b)


"\mu_k(N-100)=F\to F=0.34\cdot(90\cdot9.8-100)=266(N)"


"\Delta F=300-266=34(N)" . Answer